OpenStudy (ineedhelp10):

A rectangle and a square have the same perimeter. The square has coordinates of (-4,7); (2,7); (-4,1) and (2,1). What is the width of the rectangle if the length is 10?

OpenStudy (ineedhelp10):

@Directrix

Directrix (directrix):

@ineedhelp10 (-4,7); (2,7); (-4,1) and (2,1) I assume that these points are coordinates of consecutive vertices of the square. The distance from (-4,7)to (2,7) from (2,7) to (-4,1) from (-4,1) to (2,1) and from (2,1) to (-4,7) should be the same. You could use the distance formula to get the side of the square but if you plot the points, you will see what the side lengh has to be. Post what you get of the side of the square and then we will move to the next step, okay?

OpenStudy (ineedhelp10):

i got 6

Directrix (directrix):

Great. So, the perimeter is 6*4 = 24 for both the square and the rectange.

OpenStudy (ineedhelp10):

but its asking for the side length, so that'll mean that the side length is 6?

Directrix (directrix):

The length of the rectangle is 10. Opposite sides of a rectangle are congruent. So, the perimeter of the rectangle is 10 + 10 + w + w = 24 where w is the width of the rectange.

Directrix (directrix):

@ineedhelp10 Solve this equation for w: 10 + 10 + w + w = 24 Let me know what you get, okay?

OpenStudy (ineedhelp10):

20w^2

Directrix (directrix):

@ineedhelp10 >>but its asking for the side length, so that'll mean that the side length is 6? The side of the square is 6, yes. But, the question asks for the width of the rectangle. That is what we are working on now.

Directrix (directrix):

not 20w^2 10 + 10 + w + w = 24 Add like terms 20 + 2w = 24 Subtract 20 from both sides ---> Pick up the equation solution from here @ineedhelp10 Note: w + w = 2w. w times w = s^2

Directrix (directrix):

Here is where you are: 20 + 2w = 24 ---> Subtract 20 from both sides

OpenStudy (ineedhelp10):

w=2

Directrix (directrix):

Yes. That is what I got.

OpenStudy (ineedhelp10):

and after that?

OpenStudy (ineedhelp10):

@Directrix

Directrix (directrix):

That is it. You have answered the question. Question: What is the width of the rectangle? Answer: 2 as you worked above. @ineedhelp10